Write the equation of the line that passes through point (0, –9) with a slope of 6.
Answer by tutorcecilia(2152)

Use the slope-intercept equation of a line: y=mx+b
y=any y-value
m=the slope of the line
x=any corresponding x=value
b=the y-intercept(where the line crosses the y-axis)
.
Plug-in point (0, -9) and slope= 6 into the formula and solve for the b-term:
y=mx+b
-9=(6)(0)+b
-9=b
.
So,the equation of the line is:
y=6x-9 [you only need the values of the slope and the y-intercept].
.
checking:
=9=6(0)-9
=9=-9 [checks out]